i have one question , i would like to know , what is the rule of solving this quetion, the question is ( in one afericaon country presidential elections are held roughly every five years , if there was an election in the year 2000, when would the next four elections occur ? how do you know ? )
thansk alot

Answer by rothauserc(4718)

This is an arithmetic series,
2000, 2005, 2010, 2015, 2020
the term Xn = Xo + d(n-1) where Xo = 2000 and d = 5 (elections held every 5 years)
the next four elections are 2005, 2010, 2015, 2020
for example, the second election is
X2 = 2000 + 5(2-1)
X2 = 2000 + 5 = 2005